How many meters per minute are equal to 6 centimeters per second?
Answer by nabla(475)

(6 centimeters)* (1 meter)/(100 centimeters)=.06 meters/second
(.06 meters/second)*(60 seconds/minute)=3.6 meters/minute
check: 3.6*100=360 cm per minute. 360/60=6 cm per second.